OpenStudy (anonymous):

Determine A if: A = E1i . E2i . B

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 0 & 0 \\ 0 & -2 & 0\\0 & 0&1 \end{matrix}\right]\]

OpenStudy (anonymous):

Above is the matrixE1i

OpenStudy (anonymous):

E2i= \[\left[\begin{matrix}1 & -0.5 &0 \\ 0 & 1&0\\ 0&0&1\end{matrix}\right]\]

OpenStudy (anonymous):

and B = \[\left[\begin{matrix}3 & 2&1 \\ -2 & 0&1\\ 3&1&1\end{matrix}\right]\]